A cannon ball of mass 6kg is fired from east to west with a speed of 50km/hr at a latitudes of 30degree.whatis coriolis force
1
Expert's answer
2020-04-01T10:17:28-0400

The Coriolis force will be


"F_c=2m\Omega v\text{ sin}\lambda=\\
\space\\
=2\cdot6\cdot(7.3\cdot10^{-5})\cdot\bigg(50\cdot\frac{1000}{3600}\bigg)\cdot\text{sin}30^\circ=\\
\space\\
=6.08\cdot10^{-3}\text{ N}."

Notice that in the equation above Ω is the earth's angular velocity:


"\Omega=\frac{\Delta\phi}{\Delta t}=\frac{2\pi}{24\cdot3600}=7.27\cdot10^{-5}\text{ s}^{-1}."